OpenStudy (anonymous):

for any two events X and Y, P(X/Y)=1-P(Y/X) True or False?

OpenStudy (anonymous):

seems unlikely, we can try one and see

OpenStudy (anonymous):

what is true is that \(P(A^c)=1-P(A)\) but \[P(A|B)=\frac{P(A\cap B)}{P(B)}\] and similarly \[P(B|A)=\frac{P(A\cap B)}{P(A)}\] but one is not one minus the other

OpenStudy (anonymous):

so it's FALSE

OpenStudy (zzr0ck3r):

is this set theory or stats? im confused

OpenStudy (anonymous):

yes, it is false